OpenStudy (anonymous):

The number of bacteria, n, in a dish, after t minutes is given by n=800e^0.13t. Find the value of n when t=0. Find the rate at which n is increasing when t=15

OpenStudy (anonymous):

n when t=0 is 8800*e^0=800 rate at which it is increasing dn/dt=800*0.13t*e^0.13t=104*e^0.13t rate of increase at t=15 dn/dt=104*e^(0.13*15)=

OpenStudy (anonymous):

what is dn/dt?

OpenStudy (darkprince14):

dn/dt derivative of n with respect to t... I think?

OpenStudy (anonymous):

dn/dt is derivate of n w.r.t t which gives the rate of change of n

OpenStudy (darkprince14):

Calculus!!!

OpenStudy (anonymous):

i didnt learn derivitives yet.. lol

OpenStudy (anonymous):

Well the rate at which it is increasing will just be \(800(0.13)e^{0.13t}\). That's why exponential curves are interesting. Plug in \(t=15\).